### v2

# PLAICraft: Large-Scale Time-Aligned Vision-Speech-Action Dataset for Embodied AI

Abstract
Advances in deep generative modeling have made it increasingly plausible to train human-level embodied agents. Yet progress has been limited by the absence of large-scale, real-time, multi-modal, and socially interactive datasets that reflect the sensory-motor complexity of natural environments. To address this, we present PLAICraft, a novel data collection platform and dataset capturing multiplayer Minecraft interactions across five time-aligned modalities: video, game output audio, microphone input audio, mouse, and keyboard actions. Each modality is logged with millisecond time precision, enabling the study of synchronous, embodied behaviour in a rich, open-ended world. The dataset comprises over 10,000 hours of gameplay from more than 10,000 global participants. Alongside the dataset, we provide an evaluation suite for benchmarking model capabilities in object recognition, spatial awareness, language grounding, and long-term memory. PLAICraft opens a path toward training and evaluating agents that act fluently and purposefully in real time, paving the way for truly embodied artificial intelligence.
1 Introduction
Embodied artificial intelligence (EAI) aims to build agents that perceive, act, and learn in dynamic environments through real-time sensory-motor interaction. As Brooks [6] emphasized, such agents must “cope appropriately and in a timely fashion with changes in [their] dynamic environment” and “do something in [their] world; [they] should have some purpose in being.” These principles remain vital as research shifts toward agentic AI systems like Devin [1] and Claude Code [2] that operate autonomously across extended tasks.
Yet most modern agents are disembodied in the sense that they interact with their environment indirectly, often through high-level text interfaces and asynchronous API calls. This limits their ability to engage in temporally grounded, perceptual-motor loops that characterize embodied intelligence. While simulated environments such as Habitat [23], RoboTHOR [11], and ThreeDWorld [15] support real-time, sensorimotor learning, they often lack socially interactive or linguistically rich scenarios. Minecraft datasets like MineRL [17], OpenAI’s VPT [5], and CraftAssist [8] have driven progress in perception–action coupling and language learning. Most recently, MineDojo [14] introduced large-scale web-mined video-language data to support open-ended instruction following and knowledge transfer. However, these settings remain offline, non-interactive, heterogeneous, or lack real-time, social grounding—that is, the process of linking symbols to perceptual input, physical action, or shared environmental context.
We introduce PLAICraft, a dataset and platform designed for the study of real-time, socially situated, speech-interactive embodied agents. Built atop a modded multiplayer Minecraft server with a proximity-based voice chat plugin, PLAICraft records time-aligned streams of screen video, game output audio, player microphone input audio, keyboard, and mouse input with millisecond precision. Agents operate in persistent, open-ended worlds with other players where perception, speech, and action are tightly coupled in time. In its emphasis on rich, temporally grounded sensory-motor interaction, PLAICraft aligns with recent efforts to build datasets that support embodied learning in complex, naturalistic domains—such as autonomous driving benchmarks designed to capture diverse, real-world scenarios [31, 35].
Fig.˜1 illustrates our approach: participants play in our virtual environment and interact with other players naturally while all modalities are recorded. Unlike most prior datasets, PLAICraft emphasizes dialogue, social context, and temporal embodiment, supporting the study of skills like object permanence, reactive speech, memory, and social reasoning. Agents must act with purpose amid uncertainty, social influence, and ambiguity.
Following Brooks’ philosophy, we focus not on reducing complexity but embracing it—starting with large-scale imitation learning as a goal. Standard reinforcement learning approaches, when used in isolation, struggle in this setting. PLAICraft, like the real world, lacks explicit, global reward signals and is not structured around predefined episodic tasks. While the game mechanics do allow for death and respawn cycles, these do not constitute clean episodic resets from the agent’s perspective: the agent persists through its avatar’s deaths, and such events may even form part of emergent strategies or social play.
We have collected over 10,000 hours of data from over 10,000 players around the world. We introduce the PLAICraft dataset in Sec.˜2. The PLAICraft data collection platform is described in Sec.˜3. Sec.˜4 covers an evaluation suite designed to probe reasoning, memory, and communicative competence. We conclude in Sec.˜6 and Sec.˜8.
2 Plaicraft Dataset
We have chosen Minecraft as our environment, like many previous datasets [25, 14, 5, 8], because its rich, open-ended dynamics make learned behaviours transferable to real-world scenarios. In vanilla Minecraft, players navigate procedurally generated 3D worlds abundant with diverse terrains, resources to gather, tools to create, structures to construct, and novel discoveries to encounter. The world players interact with is also boundary-free, allowing unrestricted exploration. Minecraft provides neither explicit goals nor predefined storylines, rendering it uniquely suitable as a foundation for open-ended embodied AI research. Unlike many prior Minecraft datasets, which are mostly restricted to single-player mode with limited sets of tasks [25], limited modalities or non-systematic internet-fetched videos [14], our PLAICraft dataset is unique in the following ways.
Unbounded Multiplayer Social Environment: The open-task nature of Minecraft inherently offers an unlimited spectrum of possible tasks. Our dataset further enhances this by employing a multiplayer environment where participants are encouraged to socially interact with each other, and a proximity-based voice chat plugin for natural voice communication. This setup substantially enriches the complexity of behaviours, particularly enabling realistic social interactions. Figure 2 shows a wide range of solo and collaborative activities players have performed on the server.
Our participants comprise Minecraft enthusiasts globally, spanning diverse age groups, genders, and experience levels. Self-reported demographic statistics of participants are illustrated in Figure 3. During each gameplay session, participants receive no specific task assignments or restrictions regarding interaction areas. Instead, they are encouraged to play naturally, thinking of this not as a research project but just regular Minecraft, subject only to basic behavioural guidelines promoting mutual respect. The virtual microphone on the VM is forced to be enabled, and we encourage our players to continuously talk, especially when they play interactively with others. In each gameplay session, players have a maximum of 4 hours of play time; thus, they can perform a wide range of different activities. For example, as shown in Figure˜4, the player engaged in a wide range of interactive activities from minute 1 to minute 22, and from minute 22 to the end, the player only performed solo activities.
Consequently, our dataset exhibits extensive dynamic variability. Solo player activities span a wide spectrum, beginning with basic actions such as moving with WASD keys or breaking a dirt block and progressing to moderately complex behaviours like mining, crafting, and combating. They extend further to events that require advanced skills, including defeating the Ender Dragon, constructing intricate redstone mechanisms, or even exploiting server bugs for personal benefit. Additionally, the dataset captures numerous social interactions, ranging from simple greetings and self-introductions to collaborative activities involving team formation and cooperative gameplay. Notably, participants often form teams spontaneously, engaging in continuous interaction through collective base construction and various other activities. Players between different teams interact with each other more cautiously and less frequently. Despite rigorous moderation efforts, occasional negative interactions such as conflicts, theft, and intentional property damage are present. Such behaviours are explicitly identified, and strong measures are implemented to prevent malicious participation.
Millisecond-precision Time-aligned Multimodal Data: The Plaicraft dataset comprises tens of thousands of gameplay session recordings, each representing one session of an individual player. For every session, we simultaneously capture multiple modalities, including screen recordings, game output audio, player microphone input audio, and mouse and keyboard interactions. These modalities are recorded with millisecond-precision timestamps, facilitating precise temporal alignment across data streams. Critically, the alignment between mouse and keyboard inputs and their corresponding visual and auditory outputs is strictly causal, ensuring input actions always precede or coincide with their resultant effects. During preprocessing, mouse and keyboard data are categorized into mouse & keyboard clicks and mouse movements. Click data includes discrete "PRESS" and "RELEASE" events, whereas mouse movement data captures continuous relative positional changes.
In addition to raw multimodal data, we provide encoded data representations generated by specialized autoencoders (detailed in Section 3). Figure˜5 illustrates the temporal alignment of these encoded formats. Given the differing sampling rates of video encodings (10 frames per second) and audio encodings (75 frames per second), each video frame corresponds to 7.5 audio frames. To resolve this fractional correspondence, we define the minimal encoding unit as two consecutive video frames (200 milliseconds).
Continual global world state: Throughout the entire duration of our project, the server has maintained a single, persistent world state that has evolved continuously since its creation a year ago, without resets. Consequently, player attributes—such as position, inventories, and experience level—carry over seamlessly between their gameplay sessions, unless they die in the game, then they will lose their inventories and experience if it is not picked up soon enough. Structures built by players endure indefinitely, and any environmental changes (for example, mined blocks) become permanent. Figure 6 illustrates the development of a communal player base constructed collaboratively by roughly ten participants over a one-year period. Thus, our dataset can also be considered as the continual history of the world’s evolution from the perspectives of more than 10,000 unique players.
3 Data Collection
Gathering gameplay data often involves developing specialized platforms tailored to specific games, each requiring unique strategies for user engagement and data collection. MineRL [17], for instance, utilizes a server and plugin system that players can access for gameplay. The data set provided by MineDojo [14] is crawled from pre-existing internet platforms such as YouTube, Wikipedia, and Reddit. The PLAICraft data collection platform standarizes the data collection process, ensuring the capturing of time-aligned video, game output audio, microphone input audio, keyboard and mouse action data while providing participants the freedom to naturally and collaboratively engage within a multi-player Minecraft setting.
3.1 Platform Setup
As illustrated in Figure 7, our platform is built primarily on Amazon Web Services (AWS). The core of the architecture is a set of AWS Lambda functions that serve as the backend. These Python functions handle interactions with various AWS resources; for example, some functions read from and write to Amazon DynamoDB, the database service where we store user information and gameplay status. Each Lambda function is fronted by Amazon API Gateway, which exposes a secure endpoint for external invocation.
Users interact exclusively with the front end, which is a simple website accessible in any Chromium-based browser. The front end is hosted in Amazon S3 and delivered to users via Amazon CloudFront, a global content delivery network. Players log in with their email address or phone number and are then redirected to an Amazon EC2 instance where they can play the game.
The EC2 instance is the virtual server on which we run the Minecraft game and associated data-collection software. We pre-package an Amazon Machine Image (AMI) from an instance with all software and environment configuration. We configured the instance to immerse participants in gameplay by disabling any UI interactions other than the Minecraft window. Upon user login, a dedicated g5.2xlarge instance will be spawned from the same AMI and bound exclusively to that user. This ensures secure, isolated and uniform data recording. The Minecraft client launches in full-screen mode and cannot be minimized. We also run auxiliary monitoring scripts that track the player’s connectivity and window focus; if they detect that the player has disconnected or shifted focus away from the Minecraft window, the scripts automatically terminate the EC2 instance and upload the recorded data to S3.
The instance is equipped with NVIDIA A10G Tensor Core GPUs so that our GPU-demanding game and software can run smoothly. The instance’s display is streamed to the user through NICE DCV, a high-performance remote display protocol, at up to 60 FPS and up to 2K resolution. Our platform offers a similar experience to popular cloud gaming services such as GeForce NOW and XBOX Cloud Gaming. Moreover, since recording occur within the secure environment of the virtual machine, participant privacy is fully protected.
3.2 Minecraft Configurations
We retained most features of vanilla Java Minecraft. Players exclusively engage in survival mode with normal difficulty settings, and all three standard worlds (the Overworld, Nether, and the End) are fully accessible. To encourage social interactions, teleportation capabilities and a global chat interface are provided.
We used a combination of custom and third-party plugins to enhance the player experience. In particular, because buying unique game licenses to every participant would be impractical, we developed the Autojoin plugin, which saves and restores each player’s state (inventory, position, level, etc.) in our DynamoDB backend rather than binding it to a specific license. This decoupling allows us to reuse a limited pool of licenses across non-concurrent players, enabling the platform to scale to more than 10,000 participants.
The most critical enhancement is the integration of the Simple Voice Chat plugin, a Spigot-based modification enabling immersive, proximity-based 3D voice communication within the game. Players can communicate naturally using microphones, with audio volume dynamically adjusting based on in-game distance and relative positioning. This feature substantially enriches player interactions, elevating the dataset’s complexity and realism to a level comparable with real-world interaction datasets.
Finally, the inclusion of the Complementary Shader Pack (Unbound) significantly improves visual detail and complexity closer to real-world environments. A comprehensive list of applied plugins is provided in the Section˜A.3.
3.3 Increasing Social Interactions
Social interaction is a central focus of our dataset, and we employ multiple mechanisms to encourage social engagement. We enable a persistent multiplayer environment, integrate a voice-chat plugin, and deploy curated mods that lower friction for collaboration and communication. We further provide UI/UX features in both the web client and in-game interfaces that streamline inviting friends, forming parties, and coordinating play.
We also run live transcription using WhisperX to process players’ speech audio and their proximate hearing audio during gameplay. The resulting transcripts are analyzed by a large language model to detect and characterize meaningful conversations with other players. Based on the quality and quantity of such interactions, we allocate additional solo play time as an incentive, thereby reinforcing sustained, socially rich behavior.
3.4 Data Recording & Pre-processing
We synchronously capture all modalities on the EC2 instance using dedicated software components, as shown in Table˜1. We use the exact timestamp at which OBS’s ffmpeg muxer begins writing the video file as the global start time and align all modalities to this reference. We then apply filtering to remove noisy data and harmful content, after which the data are pre-encoded using specialized autoencoders to reduce computational burden during model training. The detailed pipeline is described in Section˜A.6 and Section˜A.7.
| Modality | Software | Raw format | Encoding method | Encoded format |
| Video | OBS | 30 FPS, 1280720, H.264 | sdxl-vae-fp16-fix [22] | 10FPS, (C=4, W=160, H=96) |
| Speaking audio | OBS | 48 kHz, stereo, AAC | Encodec 24 kHz [défossez2022highfidelityneuralaudio] | 75Hz, (D=128) |
| Hearing audio | OBS | 48 kHz, stereo, AAC | Encodec 24 kHz [défossez2022highfidelityneuralaudio] | 75Hz, (D=128) |
| Mouse movement | Forge mod | 100 Hz, (D=2) | Normalization | 100 Hz, (D=2) |
| Key presses | Forge mod | 100 Hz, (D=79) | Custom autoencoder | 50Hz, (D=16) |
3.5 Data Collection History
We do not hire experts to play since Minecraft already has the largest player number globally. With extensive media outreach and promotion efforts, we have formed a large community of voluntary players consists of Minecraft enthusiasts and experts. Figure 8 demonstrates our data collection and player base development history.
4 Evaluation
| Test Type | Spoken Instruction | Response | (ms) | (ms) | Metric |
| audio to action binding | Press A. | A pressed. | 689700 | 689850 | |
| spelling ability | Please press the letter that ends the word STOP, as in S-T-O. What letter was that? | P pressed. | 115950 | 116070 | |
| object recognition | What do you have in your right hand? | Carrot. | 548097 | 548557 | |
| world-model understanding | I’m going to shoot you with an arrow, okay? And I want you to tell me how many hearts of damage are done, okay? | Two. | 510590 | 511010 | |
| arithmetic | If you add up the blocks in the wool and granite pile, how many are there? | Five. | 1146213 | 1146573 | |
| spatial understanding | How many blocks away is the box? | Hundred. | 583714 | 584054 |
Motivated by Animal-AI [29], which adopts testing paradigms from comparative psychology; BIB [16], which assesses machines’ abilities to reason about other agents’ intentions by observing their actions; and the taxonomy of cognitive abilities outlined in Cattell-Horn-Carroll (CHC) theory [7], we publish an evaluation suite that probes how agents perceive, understand, reason, and act within PLAICraft. The idea behind this design is intuitive: if an agent truly understands what it sees and hears, it should be able to (i) talk about it, (ii) act on it, and (iii) do so within a time‑bounded, social context.
Our evaluation suite contains carefully curated prompts. Each prompt maps to one of the ten broad cognitive abilities in CHC theory [24]: fluid intelligence, quantitative knowledge, crystallized intelligence, short-term memory, visual processing, auditory processing, long-term retrieval, processing speed, decision speed, and reading/writing. Each prompt elicits a keyboard/mouse action or a concise verbal answer, enabling evaluation via exact‑match or semantic‑similarity metrics.
During evaluation, the agent’s prompt consists of its entire lived life up to the current moment, including every observation and interaction so far, and is then followed by a spoken instruction that triggers the response. A correct response from the agent may be a mouse & keyboard action ("A pressed") or a brief verbal answer ("Carrot"). The test type categorizes the cognitive ability being assessed, while the metric is the measure used to judge responses’ correctness.
Table 2 selects 7 representative prompts. The prompts probe a wide range of capabilities—including visual object recognition, world-model understanding, spatial estimation, arithmetic understanding and many others. These prompts require agents to integrate visual, audio, and action in a temporally grounded context. The full list of evaluation prompts can be found in Section˜A.13.
5 Related Work
Minecraft for Machine Learning Minecraft serves as an important tool for machine learning research [20, 34] because of its open-world nature, rich dynamics, and highly customizable engine. Consequently, numerous datasets and benchmarks have been established on Minecraft [17, 21, 14, 4]. Microsoft’s MALMO platform [18] first provided a programmable interface for RL in Minecraft; building on this, MineRL [17, 25] introduced human-demonstration datasets and benchmarks for control tasks. MineDojo [14] extended this line by aggregating web-scale gameplay videos with language, while the VPT Contractor dataset [4] contributed single-player demonstrations with aligned video and keyboard/mouse actions used to train inverse-dynamics and behavioural-cloning foundation models. The BASALT competition [25] applied human evaluation to open-ended tasks, and BEDD [21] further systematized evaluation across diverse, complex Minecraft scenarios.
EAI Datasets in Other Domains Beyond Minecraft, several large-scale datasets support embodied and interactive agents in other settings. GameNGen trains on a 900M-frame corpus of VizDoom agent self-play trajectories, providing dense video–action sequences that let a diffusion model learn a real-time game engine over long horizons [28]. Mind2Web contributes crowdsourced interaction traces over 2,350 open-ended tasks on 137 real websites, using HTML/DOM state and natural-language instructions to train and evaluate generalist web agents on real UIs rather than simulators [12]. Open X-Embodiment (OXE) aggregates over 1M real-robot trajectories from 60 datasets and 22 embodiments into a unified format with egocentric RGB, proprioception, language goals, and low-level control, enabling RT-X policies that transfer across robot platforms [32]. Waymo Open provides synchronized multi-camera images, LiDAR, and HD maps for thousands of urban and suburban driving segments, and serves as a standard benchmark for perception and motion forecasting in autonomous driving [27].
We specify and compare a few of these datasets side by side with PLAICraft in Table˜3.
EAI Evaluation Evaluating EAI agent performance in complex environments effectively requires direct comparisons between human and AI abilities. There is a growing literature on such benchmarks, including the Animal-AI Environment framework which has been used to compare human and AI performance on cognitive tasks inspired by animal psychology, providing insights into how agents reason and generalize [29]. Similarly, benchmarks such as the Baby Intuitions Benchmark (BIB) assess agents’ ability to infer goals, preferences, and the actions of others, highlighting gaps in AI’s understanding of intuitive physics and social reasoning compared to humans [16]. These approaches underscore the challenge of developing evaluation frameworks that align with human cognitive benchmarks [19].
6 Discussion
| PLAICraft | MineRL [25] | MineDojo [14] | VPT Contractor [4] | Mind2Web [12] | OXE / RT-X [32] | Waymo Open [27] | ||||
| Domain | Minecraft | Minecraft | Minecraft | Minecraft | Website GUI | Robotics | Autonomous driving | |||
| Modalities | Video; hearing&speaking audio; mouse; keyboard | Video; sim state; actions | Video + text | Video; key/mouse (IDM) | HTML/DOM; instruction | RGB(+Depth); proprio; controls | RGB; LiDAR; maps/trajectories | |||
| Scale | 10,000+ h, full modalities | 500 h, full modalities | 300,000 h video; 2.2B words. | 2,000 h, full modalities | 50 h (est. from 2,350 tasks) | 2,000 h (est. from 1M+ trajectories) | 574 h | |||
| Source | 10,000+ participants | Human demonstration (# not reported) | 730K+ YouTube videos, 6K+ Wiki and 340K+ Reddit posts | 10 paid contractors from UpWork | Crowdsourced AMT demonstrations | Pooled from 60 existing lab datasets | Waymo self-driving cars in real traffic | |||
, 100 Hz; , 100 Hz; , 48 kHz | , 20 Hz; , 20 Hz | , 20 Hz; , 20 Hz | ; ; | ||||||
, 30 Hz; , 48 kHz | , 20 Hz; , | ; | , 20 Hz | ; | ; ; | , , 10 Hz; ; | |||
| Open Task | Yes | No | No | No | No | No | No | |||
| Reset | No | Episodic | Episodic | Episodic | Episodic | Episodic | Episodic | |||
| Social | Yes | No | No | No | No | No | No |
PLAICraft enables new lines of inquiry in embodied AI research. The inclusion of speech audio and the multiplayer social environments can support agents to incorporate real-time social cues, fostering collaborative play that prior single-player datasets cannot capture. The task-free, continual world and agent states also lend themselves to continual learning paradigms, where agents must develop and maintain long-horizon memory over days or weeks of game time while continually adapting to a non-stationary environment. The same long, rich trajectories, collected from the perspective of different players, can also enable systematic study of third-person learning: training agents purely from logs of other players’ behaviour rather than from their own on-policy experience. This setting, with controllable data and compute budgets, naturally supports quantitative investigation of scaling laws for embodied agents, relating training resources to control, communication, and social coordination capabilities. The vast collection of precisely aligned high-resolution videos and ambient audios—from the diverse perspectives of thousands of players in a shared world—also makes PLAICraft well-suited for world model training, 3D, and physics simulation research.
Our data collection platform is application-agnostic: because it logs raw keyboard/mouse and audio around a fixed display stream inside a controlled VM, we can easily replace Minecraft with any desktop application (e.g., web browser, productivity or creative tools) while preserving the same data schema and time-alignment guarantees. This allows us to extend beyond Minecraft agents to general computer-use agents that can operate a full computer interface (screen, audio, keyboard, mouse) at high temporal precision. The value of computer-use agents does not stop at information workflow. Remote teleoperators of warehouse and other robots are able to actuate complex robots and other cyberphysical infrastructure entirely through a standard computer interface, albeit with the added requirement of high time-precision “dexterous” keyboard and mouse actuation. Computer use agents that learn to mimic teleoperation control inputs to embodied robot agents have the potential to become de facto control policies of such robots.
Since the initial release of a subset of our dataset, it has attracted many interests from the research community. For example, Yoo et al. [33] leveraged a 50-hour subset—collected from an anonymous player (“Alex”)—to demonstrate that a video-diffusion model can be effectively trained within a continual-learning setting on a continous stream of our dataset[30], samples are shown in Figure˜9.
7 Ethical Considerations
This research ensures the safety and privacy of participants through the guidance of our university’s office of research ethics. Study procedures are reviewed routinely by the Behavioural Research Ethics Board (BREB); and have been found acceptable on ethical grounds for research involving human subjects, as well as deemed minimal risk to participants. Participants’ data is collected with their consent, or that of their guardian, and any reasonably identifiable aspects of this data are anonymized systematically.
8 Conclusion
We introduce PLAICraft, a standardized data-collection platform and large-scale multimodal dataset that provides millisecond-aligned video, speech and hearing audio, keyboard/mouse data with rich social interactions. We also introduced an evaluation suite aimed at probing the important capabilities a human-like agent should have. PLAICraft is designed to open new avenues for research in social embodied agents, continual and third-person learning, world-model, general computer-use and teleoperation-style agents.
Supplementary Material
Appendix A Appendix
A.1 Acknowledgment
We are very grateful to Alexander Liteplo, Alex Wu, Andrew Smith, Andy Huang, Brooke Dai, Chengyuan Yao, Christopher Tardy, Daniel Crookall, David Tianyi Yin, David Yang, Edward Liang, Geo Lee, Hanson Sun, Joey Xiang, Lucas Qin, Mantaj Dhillon, Mehdi Safaee, Naveed Ghassemi, Perry Zhu, Sean Chuah, and Suzette Sun for their help in the development of the platform. We thank Hotslicer Media, Alice Xia and many others for their promotion efforts. We also appreciate the help with server moderation and community organization by David Yang, Piper, Delara, and Hiroshi (the last three are anonymous players on the server who spontaneously offered to help with moderation). In addition, we would like to send a special thanks to all the amazing players who have spent countless hours on our server, enjoying Minecraft, building a wholesome community and contributing to our research.
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada CIFAR AI Chairs Program, Inverted AI, MITACS, and Google. This research was enabled in part by technical support and computational resources provided by the Digital Research Alliance of Canada Compute Canada (alliancecan.ca), the Advanced Research Computing at the University of British Columbia (arc.ubc.ca), and Amazon.
A.2 Societal Impacts
Combining embodied AI with immersive game environments has the potential to amplify both the benefits and risks of artificial intelligence. On the upside, interactive gameplay environments like Minecraft provide a uniquely rich substrate for training and evaluating agents with social, linguistic, and perceptual competencies. Agents trained in such environments could support educational gameplay, serve as in-game tutors, or scaffold more collaborative and engaging human experiences. As embodied agents become more fluent in social cues—especially through real-time speech and interaction—they may also help simulate realistic training scenarios or serve therapeutic and accessibility functions.
At the same time, advances in speech-enabled embodied agents may blur the line between artificial and human players. Players may not always be able to distinguish between synthetic and human teammates or adversaries. While this ambiguity can enhance immersion, it may also raise concerns about deception, manipulation, or the erosion of trust. Agents that convincingly imitate social behaviours could influence players in subtle ways, especially younger users. There is a risk that such agents could be used to simulate companionship without reciprocity, or to manipulate attention and emotional responses for commercial ends.
The collection of large-scale, multimodal human gameplay data also raises important questions about consent, privacy, and downstream use. We have undergone an institutional privacy review and obtained consent for data release, and have designed our infrastructure to preserve anonymity and prevent the recording of persistent identity information. Nonetheless, the richness of multimodal recordings—particularly audio—makes full anonymization difficult. Developers and researchers using such data should exercise caution in avoiding reidentification risks and consider the broader implications of using data collected in seemingly informal or playful contexts to train powerful AI systems.
Finally, embodied AI in gaming environments may raise labour and cultural concerns. As agents become capable of performing in-game roles—e.g., assistants, performers, or economic participants—they could displace human workers in virtual economies, reshape online social norms, or saturate game environments with synthetic interactions. While our work is aimed at research and foundational exploration, we believe that careful governance, transparency, and participatory design will be essential as embodied agents become more integrated into human play and work environments.
| Label | Question |
| Interactive | Is this segment interactive? |
| Minecraft-related | Is this segment related to Minecraft gameplay? |
| Offensive | Does this segment contain offensive or inappropriate language? |
| Crafting | Are the players building or crafting something together? |
| Exploring | Are the players actively navigating the Minecraft world? |
| Teaching | Is one player helping another learn something about the game? |
| Fighting | Are the players involved in or discussing combat? |
| Game Action Quality | Is this segment a high-quality example of a gameplay interaction? |
A.3 Minecraft Configurations
We retained most features of vanilla Java Minecraft. All three worlds (the Overworld, the Nether, and the End) are accessible to players. In addition to the standard Minecraft setup, we implemented several client-side enhancements:
1.
We restricted access to in-game settings to prevent players from exiting the game; should a player leave the Minecraft window, the instance is shut down, and the data recorded up to that point is uploaded, ensuring that all captured data consists solely of gameplay.
2.
We implemented a mod to capture players’ keyboard and mouse actions with a more precise timestamp than system-level logging, bringing the recorded actions closer to the Minecraft server tick and thus better aligning cause and effect with rendered frames.
3.
We built a live monitor pipeline, using WhisperX and LLM models to monitor players’ speaking and hearing audio. We then use this as metrics to determine the quality of interaction between players and reward them based on it to encourage more meaningful and high-quality interactions.
4.
We incorporated a range of third-party client-side mods and plugins to enhance performance, visuals, and playability as shown in Table˜5.
On the server side, adjustments include:
1.
A custom mod, AutoJoin, designed to manage player dynamics. We store players’ state (inventory, level, location, etc.) in our own database. Every time a player joins, a random license will be selected and loaded. Upon joining, the plugin fetches the player’s state, overwrites the license’s state with it, and teleports the player to their last location. When a player leaves, their game state is saved back into the database, giving the player a seamless experience while allowing us to support more concurrent users.
2.
To encourage social interactions, we added a few new worlds in addition to the default three worlds: A lobby world where players can access at any time by entering the command "/minigames", this is a containerized world with boundaries where people can safely meet up with other players. From the lobby world, players can choose to go to two other worlds: a world integrated with BedWars, one of the most popular Minecraft minigames, and a world integrated with Murder Mystery, another popular minigame. Both of these game modes are added in the hope of increasing the number of participants and players’ interactions.
3.
Various third-party server-side plugins are added as shown in Table˜6:
| Name | Description |
| Complementary Shader Pack | Significantly enhances visual detail and overall graphical quality. |
| Simple Voice Chat | Proximity-based voice chat enabling real-time communication. |
| Canary | Performance optimization mod for general client-side speedups. |
| Clumps | Reduces item-entity lag by grouping dropped items into stacks. |
| EntityCulling | Skips rendering of off-screen entities to improve FPS. |
| Fastload-Reforged | Accelerates startup and world-loading times. |
| FerriteCore | Minimizes memory usage by stripping unused block states. |
| ModernFix | Addresses rendering performance issues and chunk-loading bugs. |
| Oculus | Optimizes block-update and rendering code paths for better FPS. |
| Redirector | Replaces generic implementations with faster runtime alternatives. |
| Rubidium | Speeds up core rendering operations via low-level optimizations. |
| Saturn | Bundles multiple performance tweaks for smoother gameplay. |
| SmoothBoot | Optimizes boot routines for faster game initialization. |
| Starlight | Overhauls the lighting engine for much faster chunk light updates. |
| BetterF3 | Enhanced debug overlay with customizable performance stats. |
| Cloth-Config | Provides an in-game, user-friendly configuration GUI. |
| Entity-Model-Features | Adds extra model customization hooks for entities. |
| Entity-Texture-Features | Enables per-entity texture overrides and enhancements. |
| NotEnoughAnimation | Introduces additional character animations for more realism. |
| SkinLayer3D | Renders player skin layers in true 3D for added depth. |
| Name | Description |
| AFK-Kick | Prevents idle players by kicking after inactivity. |
| BetterStructures | Adds new, interesting auto-generated structures. |
| Better-RTP | Enhances random teleportation experience. |
| ChatFilter | Filters and blocks offensive chat content. |
| Screaming Bedwars | Implements Bedwars minigame mechanics. |
| Chunky | Preloads chunks to reduce lag spikes. |
| CoreProtect | Logs and rolls back griefing actions. |
| DecentHolograms | Displays floating holograms without dependencies. |
| DiscordSRV | Bridges Discord and Minecraft chat. |
| EssentialsX | Provides core server management commands. |
| GravesX | Spawns lootable death chests at player death. |
| HideNametag | Removes player nametags for clean visuals. |
| InvisibleItemFrames | Hides item frame visuals in-game. |
| Multiverse-Core | Manages multiple worlds on one server. |
| PlaceholderAPI | Provides a uniform placeholder system for other plugins. |
| ProtocolLib | Allows manipulation of Minecraft network packets. |
| SuperVanish | Enables admins to be invisible to players. |
Collectively, these modifications create a natural, intuitive, and enjoyable environment where participants can play Minecraft as they normally would, while also providing us with standardized, time-synchronized, multi-modal gameplay data.
A.4 Additional Participants Demographics
Our participants consist of Minecraft players across the world. To lower the overall latency players will experience, we deploy our EC2 instances in various locations across the world. Figure 11 lists out our deployment regions. Players near these areas will have a low latency when playing through our streaming. Players further away from these regions are not restricted to access our platform, but their gaming experience might suffer from lagginess due to higher latency.
We do not hire specific participants to conduct our data collection process, as it is financially impractical, nor do we restrict anyone from participating. Instead, we focus on promoting our platform as a research project where participants can play free Minecraft with beautiful shaders on any pc device and make it available to everyone. We have done a series of promotions, including local poster distribution, Google Ads, YouTube and TikTok content creation. As a result, we get a constant stream of old and new participants joining our server every day. By the time of the publication, we had reached 10026 participants.
Our players come from three main sources: (1) Minecraft fans reached through our outreach. We performed an extensive media outreach and promotions, including TikTok, YouTube content creation, collaborating with famous Minecraft YouTubers, poster distributions and developed a blog website. Early growth of this category was slow, but once our promotions reached some communities, such as a school or Discord group, sign-ups surged. These players usually log the most hours and contribute the most advanced and valuable data. (2) Members of our development team. A few of our teammates who love Minecraft have added several hundred hours of play. They often help organize groups and build shared bases to boost social interactions on the server. Their contribution rate is significantly smaller than the other two groups. (3) Undergraduate students earning extra credit. In one semester we partnered with one course that offered credit for playing a set number of hours. The participation rate was high at one point, though most of these students were Minecraft beginners and did not play for long. Since the program ended, the majority of this group has stopped contributing as well.
A.5 Content Moderation
To protect dataset integrity and community health, an active moderation team oversees the game server around the clock. Supported by automated plugins and help from veteran players, moderators can promptly detect, kick, and ban misbehaving users on sight. Because every gameplay action and chat message is recorded and logged, any incidents missed in real time can be traced and addressed retrospectively. These safeguards preserve a wholesome community atmosphere, sustain growth, and limit contamination of the collected data.
A.6 Recording Software
Video & Audio Video recordings are captured using Open Broadcaster Software (OBS) running on the EC2 instance. Videos are recorded at a constant frame rate of 30 FPS with a resolution of 1280x720, encoded using the H.264 codec. OBS is specifically configured to capture only the Minecraft application window in full-screen mode, ensuring that all recorded frames pertain solely to gameplay. Audio streams are simultaneously recorded by OBS, separately capturing user microphone input and game audio output. Each audio stream is recorded at a constant sampling rate of 48 kHz in stereo format, encoded with the AAC codec. We take the exact timestamp when OBS’s ffmpeg muxer began writing the file as the start timestamp, and align all modalities to it.
Mouse & Keyboard Mouse and keyboard inputs are precisely captured using a custom-developed Forge mod integrated directly into the client-side Minecraft application. This method provides superior timing accuracy by recording when the game server actually processes input events, unlike traditional system-level logs. Input events for both mouse and keyboard are recorded on an event-driven basis with millisecond precision, maintaining a maximum polling rate of 100 Hz, resulting in a minimum interval of 10 ms between consecutive data points. Additionally, comprehensive system-level logging using pyxhook is implemented as a secondary data collection measure. These system-level logs, although less precise in timestamp accuracy, serve primarily to validate the integrity and completeness of the mod-generated input data.
A.7 Data Pre-processing
We conducted a systematic and rigorous pre-processing pipeline on the raw data, transforming it into a structured and accessible format for public release.
A.7.1 Video
Video data were extracted from the original MKV format and converted into MP4 files with a standardized frame rate of 30 FPS. We trim the start of the video to align with the player’s first keyboard clicks, and we trim off the end to match when the player quits the game. This way, we make sure the content in the video is mostly related to the actual gameplay.
A.7.2 Audio
Audio data comprised two distinct tracks, each independently extracted and stored in WAV format. The first track, representing audio captured through player microphones (referred to as audio_speak), underwent a preprocessing step using the Silero Voice Activity Detector (VAD) to isolate valid speech segments from background noise. Furthermore, speaker diarization was implemented utilizing the Pyannote model, assigning unique speaker labels to each speech segment. These speaker annotations were subsequently integrated with text transcripts generated using the Whisperx model. The second audio track (audio_hear), containing ambient audio, was processed similarly but without explicit noise filtering prior to transcription. Both audio datasets, along with their respective speaker labels, were systematically stored in a database inclusive of word-level timestamps in millisecond precision.
A.7.3 Mouse & Keyboard
Mouse data preprocessing involved segregating mouse click events from mouse movement trajectories. Mouse movement data were aggregated into fixed temporal bins of 100 ms, each represented by arrays of dimensions (2, 10), corresponding to relative x and y movements across timestamps aligned to a maximum polling frequency of 100 Hz. Each data point was assigned to the nearest bin, and intervals lacking data were zero-padded. Mouse click events, analogous in structure to keyboard event data, were similarly binned into arrays of shape (79, 10), where 79 dimensions represent 76 distinct keyboard keys in addition to the left, right, and middle mouse buttons. This consistent binning procedure facilitated unified storage of both mouse movement and combined mouse-keyboard event data in an organized database, enabling flexible data querying and analysis. See Figure 12 for details.
A.7.4 Data Encoding
To reduce computational load on GPUs during future training, all data are pre-encoded into lower-dimensional neural representations. Specifically, video data are encoded using the SDXL-VAE-FP16-Fix model, an adapted variant of the SDXL Variational Autoencoder (VAE) optimized for 16-bit floating-point precision. This encoding generates latent representations sampled at 10 FPS with dimensions of .
Audio signals, encompassing both and , are encoded utilizing the EnCodec 24 kHz model. Following prior approaches [26, 13], we employ continuous latent vectors produced by this model, characterized by dimensions with a frequency of 75Hz.
Additionally, we trained a temporal convolutional autoencoder with GRU layers to compress keyboard and mouse click event data, originally structured in arrays of shape per 100 ms interval (corresponding to one video latent frame). This autoencoder effectively compresses these inputs into a compact latent representation of . Conversely, mouse movement data, already compactly represented with dimensions per 100 ms interval, is retained without additional encoding.
A.8 Dataset Structure
The dataset is organized into multiple subfolders, each dedicated to data corresponding to an individual player. Within these player-specific directories, further subfolders contain both raw and encoded data from distinct gameplay sessions. Raw data includes video recordings in MP4 format captured at 30 FPS with a resolution of 1280x720, as well as two distinct WAV files with 48 kHz frequency representing (player microphone recordings) and (ambient recordings). An SQLite3 database contains detailed keyboard and mouse interaction data, optimized for efficient querying. The database also maintains comprehensive metadata for each session, including session identifiers, start timestamps, duration, and additional session-specific attributes.
Encoded data representations of video sessions are segmented into batches of 100 frames each and stored in PyTorch-compatible ‘.pt‘ files. Encoded audio data for both audio tracks is stored in HDF5 files, facilitating efficient access patterns, as they permit partial data retrieval without necessitating the loading of entire datasets into memory.
Additionally, a global dataset containing metadata is provided, encapsulating comprehensive details of the dataset, including session identifiers, modality availability, UNIX timestamps corresponding to recording times, and associated player metadata. This structured metadata design supports efficient and precise queries by the data loader and related software components, thereby streamlining targeted data extraction from the extensive dataset.
A.9 Data Versioning
There are 4 versions of the dataset, version 5 to version 8. You can find the version of each session in the info.txt under each session directory. Versions before 5 were used for testing purposes and thus not included in the dataset (the world is reset after the test). Versions 5-7 are versions of the data we recorded during the test launch period, which count towards less than 5% of the entire dataset. They have minor issues in their audio modalities and we do not recommend using them for audio training. These issues are fixed after version 8. Here is the detailed changelog of the versioning:
| Version | Change History | Known Issues |
| 5 | Initial test launch with all modalities | Audio has unexpected artifacts; keys “enter”, “esc”, and “backspace” not recorded in chat window (fixed in released version) |
| 6 | Fixed the keylogger issue in version 5 | Audio has unexpected artifacts; bad FPS occasionally |
| 7 | Added F1, F3, F5 keys; removed key binds for P, R, K (related to shaders config); added optimization mods that boost FPS | Audio has unexpected artifacts |
| 8 | Fixed the audio-artifacts issue | None |
A.10 Data Unit Format
Our dataloader functions with a global metadata database, which contains game sessions metadata like who played the session, what modalities are available in these recordings, the start time of each gameplay session, etc. The dataloader can then load the data from any specified sessions or players or from the entire database. As mentioned previously, due to the frequency discrepancy between video and audio, we treat 2 video latent frames, i.e. a 200ms window, as our minimal unit. Thus the dataloader will return a dict object as illustrated in Table˜8.
| Key | Shape |
| Metadata | [B] |
| Video tensor | [B, T, 2, 4, 96, 160] |
| audio_speak | [B, T, 15, 128] |
| audio_hear | [B, T, 15, 128] |
| Mouse movement | [B, T, 2, 10, 2] |
| Key press | [B, T, 2, 5, 16] |
[…]
A.11 Automatic Data Annotation
To identify subsets of highly interactive data, particularly relevant for Embodied AI, we developed a methodology leveraging audio transcribed by WhisperX [3], and the capabilities of large language models (LLMs). We first defined highly interactive data as continuous, game-related interactions between at least two players, within a specific time window. Notably, a key indicator of interactivity was the presence of rich, conversational audio. We used WhisperX-generated audio transcripts to create a dataset and identify these interactions. To preserve conversational coherence, we split the transcript into segments such that a new segment begins after a pause of at least 10 seconds between utterances. These segments represent potential areas of interactivity and Minecraft social interactions. Given the scale of the resulting dataset, manual annotation of segments was infeasible. Instead, we leveraged recent advances in large language models to perform automated annotation. A binary classification approach using a 4-shot prompt demonstrated high accuracy and efficiency on a validation subset. Based on this evaluation and overall cost considerations, we used models such as DeepSeek-V3 [10] and DeepSeek-R1 [9] to label the full dataset with binary annotations. Table 4 demonstrates the labeling scheme. Segments assigned specific labels were then identified as highly interactive and selected for further analysis. Figure˜10 showcases the interaction portion of a few randomly-selected players (the names are the in-game assigned names) and the weekly accumulation of their highly interactive hours.
A.12 Additional Visual Examples
We demonstrate more visual examples in Fig.˜13, Fig.˜14, Fig.˜15. The visualization follows the same format as Fig.˜2, but here we also depict temporal progression: frames within each row correspond to a single example, are spaced 1 second apart, and progress from left to right in time.
A.13 Complete Evaluation Prompt Table
| ID | Test Type | Spoken Instruction | Response | Metric | CHC | Video | Rstart (ms) | Rend (ms) |
| 1 | audio to action binding | Please press Space. | Space pressed. | 1 | 67970 | 68120 | ||
| 2 | audio to action binding | Please press Shift. | Shift pressed. | 1 | 73120 | 73650 | ||
| 3 | audio to action binding | Please press T. | T pressed. | 1 | 89240 | 89410 | ||
| 4 | spelling ability | Please press the letter that ends the word STOP, as in S-T-O… | P pressed. | 1 | 115950 | 116070 | ||
| 5 | spelling ability | Please press the letter that ends big, as in B-I… | G pressed. | 1 | 136930 | 137050 | ||
| 6 | audio to action binding | Please press the letter Q. | Q pressed. | 1 | 161850 | 162020 | ||
| 7 | action recognition | What did you just close? | Door. | 1 | 200116 | 200550 | ||
| 8 | internal state recognition | What is that green dot? | XP. | 1 | 435183 | 435866 | ||
| 9 | world model understanding | I’m going to shoot you with an arrow, okay? And I want you to tell me how many hearts of damage are done, okay? | Two. | 1 | 482750 | 483500 | ||
| 10 | object recognition | What do you have in your right hand? | Carrot. | 1 | 510550 | 511016 | ||
| 11 | object recognition | What do you have in your left hand? | Torch. | 1 | 514683 | 515383 | ||
| 12 | color understanding | What color sheep am I standing next to? | Black. | 1 | 534633 | 535216 | ||
| 13 | color understanding | What color sheep am I standing next to? | Gray. | 1 | 548100 | 548566 | ||
| 14 | color understanding | What color flower am I standing next to and looking at? | White. | 1 | 558866 | 559500 | ||
| 15 | color understanding | What color of flower am I looking at? | Red. | 1 | 566966 | 567433 | ||
| 16 | color understanding | What color of flower am I looking at? | Yellow. | 1 | 572383 | 572899 | ||
| 17 | type recognition | What is the stack of stuff that I’m standing next to made out of? | Wool. | 1 | 1111099 | 1111616 | ||
| 18 | arithmetic | If you add up the number of blocks in the wool pile and the number of blocks in the granite pile, how many blocks total are there? | Five. | 1 | 1182683 | 1183299 | ||
| 19 | arithmetic | Now, if you add up the number of blocks in the granite pile and the wool pile, how many blocks are there total? | Four. | 1 | 1199099 | 1199599 | ||
| 20 | arithmetic | I am going to place a block in the granite pile. How many blocks are in the granite pile now? | Three. | 1 | 1235883 | 1236333 | ||
| 21 | object constancy | Now, jump up. Can you see me? | Yes. | 2 | 605866 | 606200 | ||
| 22 | object constancy | Jump up. Did I move? | No. | 2 | 615450 | 615899 | ||
| 23 | audio to action binding | Press W. | W pressed. | 2 | 679650 | 679740 | ||
| 24 | audio to action binding | Press S. | S pressed. | 2 | 684010 | 684390 |
| ID | Test Type | Spoken Instruction | Response | Metric | CHC | Video | Rstart (ms) | Rend (ms) |
| 25 | audio to action binding | Press A. | A pressed. | 2 | 689700 | 689850 | ||
| 26 | audio to action binding | Press D. | D pressed. | 2 | 692750 | 692940 | ||
| 27 | audio to action binding | Press S. | S pressed. | 2 | 697070 | 697150 | ||
| 28 | audio to action binding | Do it again. Press S. | S pressed. | 2 | 709500 | 709700 | ||
| 29 | audio to action binding | Press W. | W pressed. | 2 | 713210 | 713440 | ||
| 30 | working memory, object constancy | Have you been here before? | Yes. | 2 | 1456133 | 1456466 | ||
| 31 | working memory, object constancy | Will we be someplace we’ve been before or will we be someplace new? So say old or new. | Old. | 2 | 1796900 | 1797250 | ||
| 32 | object recognition | Can you see a spider? | Yes. | 2 | 1852349 | 1852683 | ||
| 33 | object recognition | Can you see any sheep? | Yes. | 2 | 1857016 | 1857549 | ||
| 34 | type recognition | What kind of block am I standing on? | Sand. | 2 | 1926616 | 1927033 | ||
| 35 | object recognition | What am I in right now? | Water. | 2 | 1940433 | 1940783 | ||
| 36 | world model understanding | Is the sheep dead? | No. | 3 | 103183 | 103466 | ||
| 37 | world model understanding | Is the sheep dead? | Yes. | 3 | 110816 | 111266 | ||
| 38 | symbol recognition, numerosity | How many pieces of wool are in the box? | Four. | 3 | 319983 | 320266 | ||
| 39 | symbol recognition, numerosity | How many different kinds of things are in the box? | Two. | 3 | 324983 | 325166 | ||
| 40 | symbol recognition, numerosity | How many pieces of wood are in the box? | Twelve. | 3 | 346466 | 346983 | ||
| 41 | type recognition | What is the first thing that I put in the box? | Wood. | 3 | 487899 | 488516 | ||
| 42 | spatial understanding | Approximately how many blocks away is the box? | Hundred. | 3 | 583549 | 584283 | ||
| 43 | object recognition | Morgan, in front of me, what is that? | Zombie. | 3 | 849333 | 849750 | ||
| 44 | object recognition | Morgan, are those torches? | Yeah. | 3 | 895549 | 896200 | ||
| 45 | symbol recognition, numerosity | How many pieces of wood? | Six. | 3 | 1669933 | 1670616 | ||
| 46 | reading | What is the name of the last person who left the game? | Brent. | 3 | 1908099 | 1908866 | ||
| 47 | reading | Okay, what is the name of the person who joined just after me? | Chloe. | 3 | 1942033 | 1942466 | ||
| 48 | long term memory | What is the middle item in the chest? | Carrot. | 3 | 2174783 | 2175283 |
A.14 Evaluation Prompt Visual Examples
We pick a few prompts and show the nearest frames before response for visualization in Table 11 and 12.
| Nearest Frame | Spoken Instruction | Response |
What do you have in your right hand? | Carrot. |
What do you have in your left hand? | Torch. |
What color sheep am I standing next to? | Gray. |
What color of flower am I looking at? | Red. |
What is the stack of stuff that I’m standing next to made out of? | Wool. |
| Nearest Frame | Spoken Instruction | Response |
I am going to place a block in the granite pile. How many blocks are in the granite pile now? | Three. |
Now, jump up. Can you see me? | Yes. |
Can you see a spider? | Yes. |
Can you see any sheep? | Yes. |
What am I in right now? | Water. |